OpenStudy (anonymous):

Gabriel is making a mixture of compost and soil to use for a special plant. He wants his final mixture to be 2 parts compost to 10 parts potting soil. He wants to end up with 6 kilograms of mix. How many kilograms of compost should he use

OpenStudy (anonymous):

Think about it in terms of an equation... 2parts + 10parts = 6kg. how much is one "part"?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

if one part is 16 kilograms, then: (2)(16)kg + (10)(16)kg = 192kg

OpenStudy (anonymous):

solve it like an algebra problem, 2P + 10P = 6, P = ?

OpenStudy (anonymous):

I'm lost. 10+6= 16p? Then?

OpenStudy (anonymous):

Here's how I would do it: 2P + 10P = 6 12P = 6 12P/12 = 6/12 P = 1/2 so one part is one half of a kilogram

OpenStudy (anonymous):

he's using 2 parts compost. figuring out how much compost he's going to use it pretty easy now

OpenStudy (anonymous):

so 8?

OpenStudy (anonymous):

one part is 0.5kg, so 2 parts is 1kg

OpenStudy (anonymous):

and it makes sense if you plug in these numbers in the original questions.... if he uses 1kg compost, and 5kg dirt, he'll end up with 6kg soil